The concentration of Cd in lake sediment digestate is 3 µg/L. If 5 g of dry sediment was used, and the total digestate volume is 25 mL, what is the Cd concentration in sediment?

Wouldn't the mass of Cd be the 5g?
Answer #1

3 µg/L means  3 µg Cd remain in 1 L lake water.

5 g of dry sediment was used, and the total digestate volume is 25 mL = 0.025 L

thus

mass of Cd in 0.025 L lake water =  3 µg/L * 0.025 L = 0.075 µg = 0.075 * 10^-6 g

and

Cd concentration in sediment = 0.075 * 10^-6 g * 10^6 / 5 = 0.015 ppm (answer)